The fence around my rectangular backyard is 46 feet long. My yard is 2 feet longer than twice the width. What is the width of my yard? What is the length?
Answer by Cromlix(4381)

Width = x
Length = 2 + 2x
Perimeter = 2 x length + 2 x width
46 = 2 x (2 + 2x) + 2(x)
46 = 4 + 4x + 2x
46 = 4 + 6x
46 - 4 = 6x
42 = 6x
7 = x
Width = 7ft
Length = 16ft
Hope this helps.
:-)